what is y=4x-3
: what is y=4x-3

Answer by elima(1423)

Do you need to graph?
If so, you choose a value for x and solve for y;
y=4x-3
x=1
y=4(1)-3
y=1
-----------
x=0
y=4(0)-3
y=-3
-------------
x=2
y=4(2)-3
y=5
:)
--------------
graph (300,300, -5,5,-5,8, 4x-3)